### v1

Abstract
Context. Theories of the formation and evolution of small bodies in planetary systems predict that they may escape into interstellar space at any time, leaked from extrasolar Oort clouds or ejected following close encounters with local planets or their host stars. After having characterized just two such interlopers —1I/2017 U1 (‘Oumuamua) and 2I/Borisov— more questions were raised than answered. Assessing the recently discovered interstellar comet 3I/ATLAS will only broaden our understanding of this complex topic.
Aims. Here, we investigate the spectral, cometary, and rotational properties of 3I/ATLAS as well as its dynamical context.
Methods. We identified the spectral type of 3I/ATLAS from the visible reflectance spectrum and used photometric observations to derive its level of activity and rotational properties. Observational data were obtained with the OSIRIS camera spectrograph at the 10.4 m Gran Telescopio Canarias and the Two-meter Twin Telescope. We used -body simulations and statistical analyses of Gaia DR3 data to investigate the origin of 3I/ATLAS and its Galactic background.
Results. Interstellar comet 3I/ATLAS has a visible spectrum consistent with that of a D-type asteroid, and has a spectral slope =14.60.2 %/1000 Å in the 5000–9000 Å range, which is similar to the one of ‘Oumuamua but redder than that of 2I/Borisov and most solar system comets. It has a conspicuous coma and its rotation period is 16.790.23 h. The heliocentric components of its Galactic velocity were km s-1 with a radiant in Sagittarius (2950500008, 19069300012). The analysis of a sample of kinematic analogs of 3I/ATLAS extracted from Gaia DR3 suggests that its parent system is part of the Galactic thin disk and includes a solar-like star with slightly sub-solar metallicity.
Conclusions. The results from the physical characterization of 3I/ATLAS further support the idea that extrasolar debris is not too different from the one found in the solar system and it is the result of similar formation processes.
Key Words.:
comets: individual: C/2025 N1 (ATLAS) – comets: general – techniques: spectroscopic – techniques: photometric – methods: numerical – celestial mechanics1 Introduction
Interstellar objects are like cosmic bottled messages that convey information about the planetary systems where they formed for anyone to analyze, connecting distant worlds meant to never interact. Any debris in the form of asteroids and comets that escapes a mature planetary system will move with values of its heliocentric Galactic velocity components similar to those of its parent star. In the absence of significant scattering due to close encounters with other stars, this kinematic signature will remain relatively unaltered for some time. At least two comets — C/1980 E1 (Bowell) and C/2024 L5 (ATLAS) — have been recently ejected from the solar system after experiencing planetary encounters at close range (de la Fuente Marcos et al., 2024). In contrast, we had two visitors from beyond the solar system — 1I/2017 U1 (‘Oumuamua) and 2I/Borisov — and they left more questions raised than answered (see, e.g., Williams 2017; Meech et al. 2017; Micheli et al. 2018; de León et al. 2019, 2020; Deam et al. 2025). The study of more of these visitors from beyond the solar system will only broaden our understanding of this subject.
Object A11pl3Z was initially reported on July 1, 2025, by the Asteroid Terrestrial-impact Last Alert System (ATLAS, Tonry et al. 2018) observing from Rio Hurtado in Chile. It was announced as the third interstellar object on July 2 (Collaboration et al., 2025b) and designated 3I/ATLAS = C/2025 N1 (ATLAS). Follow-up and precovery observations (Ye et al., 2025b, a; Valdes et al., 2025; Collaboration et al., 2025a) revealed a highly eccentric, hyperbolic orbit (eccentricity 6). Here, we investigate observationally the spectral properties of 3I/ATLAS in the visible region using low-resolution spectroscopy, its cometary activity using high SNR images, and its rotational state from photometric observations. Our goal is to place 3I/ATLAS within the context of both known interstellar objects and solar system comets. In addition, we use its orbital solution to explore numerically its dynamical evolution. This Letter is organized as follows. In Sect. 2, we analyze the observed properties of 3I/ATLAS, comparing with those of known interstellar objects and solar system comets. Our results are discussed in Sect. 3 and our conclusions are summarized in Sect. 4. Observations, data reduction, spectral extraction, and other technical details and input data are included as appendices.
2 Results
Here, we use image enhancement algorithms applied to high-SNR images, reflectance spectroscopy, and photometric observations to investigate the physical properties and rotational state of 3I/ATLAS. The dynamical status of this comet was studied using -body simulations and its Galactic context is explored applying data mining techniques. Observational details are included in Appendices A, B, C, and D. Appendix E provides information on the numerical integrations.
2.1 Cometary activity
We generated a deep, stacked image by co-adding 227 exposures (a total of 3.15 h) from the TTT telescope obtained on the night of 2 July, that is shown in Fig. 1. Overlaid contour delineates the coma boundary at the level above the background sky. At this threshold, the surface brightness of the edge is mag arcsec-2 in the AB magnitude system. Details of the reduction and photometric calibration are given in Appendix B. The comet shows a well‐defined coma, extending over a projected area of . The major axis of this ellipse-like region is oriented at a phase angle of . At the time of observation, the geocentric distance was 3.44 au, so the physical dimensions of approximately 26 400 km 24 700 km.
2.2 Spectral properties
We studied the presence of typical cometary emission lines/bands in cometary comas (see. e.g.. Cochran et al. 2012). To do this, we proceeded as in de León et al. (2020), we obtained and analyzed the continuum-subtracted spectrum of 3I/ATLAS (see Fig. 5). We first noticed that no cometary gas emission was detected as also reported by Alvarez-Candal et al. (2025) and Opitom et al. (2025), thus the observed coma is mainly made of dust. Then we derived an upper limit for the most prominent gas emission band of comets in this spectral range, the CN (0–0) at 3880 Å, and obtained mol s-1 using the method described in Appendix C. Our results are consistent with the value reported in Alvarez-Candal et al. (2025). Ours is larger because the of their spectrum in the CN band spectral range is larger as a result of their much higher spectral resolution. Notice that measured production rates mol s-1 are unusual (A’Hearn et al., 1995) so 3I/ATLAS looks like a normal comet at this solar distance.
On the other hand, the normalized reflectance spectrum of interstellar comet 3I/ATLAS is compatible with a D-type classification, following DeMeo et al. (2009), and similar to the red spectra of the trans-Neptunian objects and Centaurs. We computed its reflectivity gradient by using a simple linear fit to its reflectance spectrum, with the same wavelength intervals and normalization wavelength as in Opitom et al. (2025) for comparison purposes: (18.50.5)%/1000 Å in the 5000-7000 Å range, (11.60.4)%/1000 Å in the 7000-9000 Å range, and (14.60.2)%/1000 Å in the 5000-9000 Å range.
2.3 Rotational state
Time-series photometry is presented in Fig. 3, which includes observations from three nights (Table 1). Data with excessively high noise due to poor seeing were removed from the plot, but were taken into account during the normalization of each light curve by subtracting the nightly median magnitude, which allowed them to be combined. The light curve shows a low amplitude of 0.2 magnitudes, which is typical for an object with cometary activity where the variability in light reflection from the nucleus’s rotation is attenuated by the presence of a surrounding homogeneous coma (Licandro et al., 2000). This is also consistent with observations in other preliminary works on 3I/ATLAS, such as Seligman et al. (2025), which, although they did not find a periodicity, they indicate that the brightness variations are small. A clear trend is observed in the TTT data from each night of observation. We estimate the rotation period of 3I/ATLAS to be h. Further details are provided in Appendix D.
2.4 Dynamics
The kinematic properties of interstellar interlopers are expected to preserve for the most part the kinematic signatures of their host stars. Analyzing the pre-encounter trajectory of 3I/ATLAS might shed some light on its true origin. We performed integrations backward in time of 1000 control orbits of 3I/ATLAS generated by the Monte Carlo using the Covariance Matrix (MCCM) method described by de la Fuente Marcos & de la Fuente Marcos (2015) and modified here to work with hyperbolic orbits. A statistical analysis (median and 16th and 84th percentiles) of the results indicates that at pc from the Sun and 3104 yr into the past, 3I/ATLAS was moving inwards, at 58.000.02 km s-1 — 1I/2017 U1 (‘Oumuamua) had km s-1, e.g. Mamajek (2017) and C/2019 Q4 (Borisov) had km s-1 (de León et al., 2020) — and projected towards (radiant or antapex) , (2950500008, ) in the constellation of Sagittarius with Galactic coordinates , , and ecliptic coordinates , . The heliocentric components of its Galactic velocity were km s-1. These values were computed as described by de la Fuente Marcos & de la Fuente Marcos (2019). The components of the solar-motion corrected heliocentric Cartesian velocities of 3I/ATLAS are (62.40.8, 7.20.5, 26.20.4) km s-1. We have searched Gaia third data release (DR3, Gaia Collaboration et al. 2016, 2023) looking for kinematic analogues of 3I/ATLAS towards the region of the radiant. Figure 4 shows 3I/ATLAS in its Galactic kinematic context.
Answering the question of where 3I/ATLAS is headed next after its fly-by with the Sun requires the analysis of a similar set of direct -body simulations (also with MCCM), but forward in time. At pc from the Sun and 3104 yr into the future, this object will be receding from us at 58.010.02 km s-1 towards (apex) , (9526003, ) in the constellation of Gemini with Galactic coordinates , , and ecliptic coordinates , . The heliocentric components of its Galactic velocity will be km s-1.
3 Discussion
Our spectroscopic results are in agreement with the ones reported by Opitom et al. (2025), and are of the order of those reported in Alvarez-Candal et al. (2025). In addition, Seligman et al. (2025) reported a slope of 18%/1000 Å from their ’,’,’,’ colors and (17.10.2)%/1000 Å from their spectrum, consistent with the values provided in this work. The spectral slope of 3I/ATLAS is redder than the nominal values adopted by Jewitt & Seligman (2023) for 1I/’Oumuamua (155 %/1000 Å) and 2I/Borisov (121 %/1000 Å).
Considering Nissen (2004), fig. 1.3, the components of the solar-motion corrected heliocentric Cartesian velocities of 3I/ATLAS place it in the thin disk as the value of is 67.68 km s-1 and =7.2 km s-1. The Toomre diagram indicates that the thin disk is inside 85 km s-1 with the thick disk in the interval (85, 180) km s-1 and the halo outside 180 km s-1. In addition to this kinematic constraint, the analysis of a sample of kinematic analogues of 3I/ATLAS from Gaia DR3 gives a value of [Fe/H]=0.040.14 (see Appendix F), which is somewhat sub-solar, but far from 1.2[Fe/H]0.4 in the thick disk (see, e.g., Nissen 2004). An origin in the thick disk is favored by Hopkins et al. (2025).
4 Conclusions
In this Letter, we presented observations of interstellar comet 3I/ATLAS obtained with GTC and TTT. The former were used to derive its spectral class and visible slope; the latter provided the rotational period and insights on its cometary activity. Direct -body simulations and statistical analyses were carried out to investigate the context of its orbital evolution. Our conclusions can be summarized as follows.
1.
High SNR images clearly show that 3I/ATLAS is active at 3.44 au , with a well‐defined dust coma that extends over a projected area of ( km km at the comet geocentric distance).
2.
3I/ATLAS has a visible reflectance spectrum somewhat consistent with that of Solar System comets and D-type asteroidd, with a spectral slope of (18.50.5)%/1000 Å in the 5000–7000Å range, (11.60.4)%/1000 Å in the 7000–9000Å range, and (14.60.2)%/1000 Å in the 5000–9000 Å range.
3.
No signatures of gas emission is observed in 3I/ATLAS visible spectrum, which is usual in solar system comets observed at au. We obtained an upper limit for the CN production rate mol s-1.
4.
The 3I/ATLAS light curve in the band exhibits a low-amplitude variation with a period of 16.790.23 h.
5.
The analysis of an extensive set of -body simulations indicate that 3I/ATLAS came from interstellar space and it will return back to it after reaching its perihelion on 2025 October 29. Its kinematic signature is consistent with an origin in the Galactic thin disk.
Follow-up observations of this object will provide further information on the properties of its parent planetary system.
Acknowledgements.
RdlFM acknowledges funding from the “ENIMUS” Advanced Grant from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement ID 101097905). This work was partially supported by the Spanish ‘Agencia Estatal de Investigación (Ministerio de Ciencia e Innovación)’ under grant PID2020-116726RB-I00 /AEI/10.13039/501100011033. This article is based on observations made in the Two-meter Twin Telescope (TTT111http://ttt.iac.es) sited at the Teide Observatory of the Instituto de Astrofísica de Canarias (IAC), that Light Bridges operates in Tenerife, Canary Islands (Spain). The Observing Time Rights (DTO) used for this research were consumed in the PEI “PLANETIX25”. This research used storage and computing capacity in ASTRO POC’s EDGE computing center at Tenerife under the form of Indefeasible Computer Rights (ICR). The ICR were consumed in the PEI ”PLANETIX25” with the collaboration of Hewlett Packard Enterprise and VAST DATA. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.References
- A’Hearn et al. (1995) A’Hearn, M. F., Millis, R. C., Schleicher, D. O., Osip, D. J., & Birch, P. V. 1995, Icarus, 118, 223
- Akhlaghi & Ichikawa (2015) Akhlaghi, M. & Ichikawa, T. 2015, ApJS, 220, 1
- Alvarez-Candal et al. (2025) Alvarez-Candal, A., Rizos, J. L., Lara, L. M., et al. 2025, arXiv e-prints, arXiv:2507.07312
- Bertin (2006) Bertin, E. 2006, in Astronomical Society of the Pacific Conference Series, Vol. 351, Astronomical Data Analysis Software and Systems XV, ed. C. Gabriel, C. Arviset, D. Ponz, & S. Enrique, 112
- Bertin et al. (2002) Bertin, E., Mellier, Y., Radovich, M., et al. 2002, in Astronomical Society of the Pacific Conference Series, Vol. 281, Astronomical Data Analysis Software and Systems XI, ed. D. A. Bohlender, D. Durand, & T. H. Handley, 228
- Bohlin et al. (2014) Bohlin, R. C., Gordon, K. D., & Tremblay, P. E. 2014, PASP, 126, 711
- Cepa (2010) Cepa, J. 2010, in Astrophysics and Space Science Proceedings, Vol. 14, Highlights of Spanish Astrophysics V, ed. J. M. Diego, L. J. Goicoechea, J. I. González-Serrano, & J. Gorgas, 15
- Cepa et al. (2000) Cepa, J., Aguiar, M., Escalera, V. G., et al. 2000, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 4008, Optical and IR Telescope Instrumentation and Detectors, ed. M. Iye & A. F. Moorwood, 623–631
- Chambers et al. (2016) Chambers, K. C., Magnier, E. A., Metcalfe, N., et al. 2016, arXiv e-prints, arXiv:1612.05560
- Cochran et al. (2012) Cochran, A. L., Barker, E. S., & Gray, C. L. 2012, Icarus, 218, 144
- Collaboration et al. (2025a) Collaboration, Z. T. F., Ye, Q. Z., Nastasi, A., et al. 2025a, Minor Planet Electronic Circulars, 2025-N51
- Collaboration et al. (2025b) Collaboration, Z. T. F., Ye, Q. Z., Robinson, J., et al. 2025b, Minor Planet Electronic Circulars, 2025-N12
- de la Fuente Marcos & de la Fuente Marcos (2015) de la Fuente Marcos, C. & de la Fuente Marcos, R. 2015, MNRAS, 453, 1288
- de la Fuente Marcos & de la Fuente Marcos (2019) de la Fuente Marcos, R. & de la Fuente Marcos, C. 2019, A&A, 627, A104
- de la Fuente Marcos et al. (2024) de la Fuente Marcos, R., de la Fuente Marcos, C., & Aarseth, S. J. 2024, A&A, 690, A395
- de León et al. (2020) de León, J., Licandro, J., de la Fuente Marcos, C., et al. 2020, MNRAS, 495, 2053
- de León et al. (2019) de León, J., Licandro, J., Serra-Ricart, M., et al. 2019, Research Notes of the American Astronomical Society, 3, 131
- Deam et al. (2025) Deam, S. E., Bannister, M. T., Opitom, C., et al. 2025, arXiv e-prints, arXiv:2507.05051
- DeMeo et al. (2009) DeMeo, F. E., Binzel, R. P., Slivan, S. M., & Bus, S. J. 2009, Icarus, 202, 160
- Gaia Collaboration et al. (2016) Gaia Collaboration, Prusti, T., de Bruijne, J. H. J., et al. 2016, A&A, 595, A1
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
- Giorgini (2011) Giorgini, J. 2011, in Journées Systèmes de Référence Spatio-temporels 2010, ed. N. Capitaine, 87–87
- Giorgini (2015) Giorgini, J. D. 2015, in IAU General Assembly, Vol. 29, 2256293
- Hopkins et al. (2025) Hopkins, M. J., Dorsey, R. C., Forbes, J. C., et al. 2025, arXiv e-prints, arXiv:2507.05318
- Jewitt & Seligman (2023) Jewitt, D. & Seligman, D. Z. 2023, ARA&A, 61, 197
- Landolt (1992) Landolt, A. U. 1992, AJ, 104, 340
- Lang et al. (2010) Lang, D., Hogg, D. W., Mierle, K., Blanton, M., & Roweis, S. 2010, AJ, 139, 1782
- Licandro et al. (2000) Licandro, J., Serra-Ricart, M., Oscoz, A., Casas, R., & Osip, D. 2000, AJ, 119, 3133
- Mamajek (2017) Mamajek, E. 2017, Research Notes of the American Astronomical Society, 1, 21
- Meech et al. (2017) Meech, K. J., Weryk, R., Micheli, M., et al. 2017, Nature, 552, 378
- Micheli et al. (2018) Micheli, M., Farnocchia, D., Meech, K. J., et al. 2018, Nature, 559, 223
- Nissen (2004) Nissen, P. E. 2004, in Origin and Evolution of the Elements, ed. A. McWilliam & M. Rauch, 154
- Opitom et al. (2025) Opitom, C., Snodgrass, C., Jehin, E., et al. 2025, arXiv e-prints, arXiv:2507.05226
- Popescu et al. (2025) Popescu, M. M., Tatsumi, E., Licandro, J., et al. 2025, PSJ, 6, 42
- Press & Rybicki (1989) Press, W. H. & Rybicki, G. B. 1989, ApJ, 338, 277
- Schleicher (2010) Schleicher, D. G. 2010, AJ, 140, 973
- Seligman et al. (2025) Seligman, D. Z., Micheli, M., Farnocchia, D., et al. 2025, arXiv e-prints, arXiv:2507.02757
- Tonry et al. (2018) Tonry, J. L., Denneau, L., Heinze, A. N., et al. 2018, PASP, 130, 064505
- Valdes et al. (2025) Valdes, F., :unkn, Valvasori, A., et al. 2025, Minor Planet Electronic Circulars, 2025-N44
- Williams (2017) Williams, G. V. 2017, Minor Planet Electronic Circulars, 2017-V17
- Ye et al. (2025a) Ye, Q. Z., Collaboration, Z. T. F., Abdelaziz, A. M., et al. 2025a, Minor Planet Electronic Circulars, 2025-N35
- Ye et al. (2025b) Ye, Q. Z., Collaboration, Z. T. F., Lee, H. J., et al. 2025b, Minor Planet Electronic Circulars, 2025-N22
Appendix A Observations and Data Reduction
Observations of interstellar comet 3I/ATLAS were performed soon after its discovery and confirmation of interstellar nature, using the telescope facilities located at the Canarian Observatories (OCAN), managed by the Instituto de Astrofísica de Canarias (Spain). Observational circumstances are described in the following sections and summarized in Table 1.
| Observing date | N. images | UTstart | UTend | Telescope | Instrument | Filter | (s) | (au) |
| 2025-07-04 | 3 | 00:24 | 00:40 | GTC | OSIRIS+ | R300R | 300 | 4.42 |
| 2025-07-04 | 2 | 00:41 | 00:52 | GTC | OSIRIS+ | R300B | 300 | 4.41 |
| 2025-07-02 | 326 | 21:27 | 02:46 | TTT | Andor Ikon-L | ’ | 50 | 4.44 |
| 2025-07-03 | 232 | 21:46 | 01:42 | TTT | Andor Ikon-L | ’ | 45 | 4.41 |
| 2025-07-05 | 307 | 21:29 | 03:12 | TTT | Andor Ikon-L | ’ | 50 | 4.34 |
A.1 Spectroscopy
Spectroscopic observations of 3I/ATLAS in the near-UV and visible wavelengths (3500–9200 Å) were obtained using OSIRIS+ camera-spectrograph (Cepa et al., 2000; Cepa, 2010) at the 10.4m Gran Telescopio Canarias (GTC), located at the El Roque de Los Muchachos Observatory (La Palma, Spain). The OSIRIS+ instrument was upgraded in January 2023 to have enhanced sensitivity in the blue. It is equipped with 2k4k pixel detector that provides a 7.87.8 arcmin2. We used the R300B (3500–7500 Å, dispersion of 4.96 Å pixel-1 for a 0.6” slit) and R300R (4800–9200 Å, dispersion of 7.74 Å pixel-1 for a 0.6” slit) grisms and a 1.5” slit width oriented in the parallactic angle. The telescope tracking was at the comet’s proper motion. Two spectrophotometric standard stars (Feige 110 and Ross 640) were observed to flux calibrate the spectra. In the same way, to obtain the relative reflectance spectrum of the comet, we also observed two solar analog stars — SA110-361 and SA112-1333 — from the Landolt catalog (Landolt, 1992).
Data reduction included bias and flat-field correction, sky background subtraction and one-dimensional aperture extraction, corresponding to the pixel where the peak intensity of the cometary profile decayed to its 10% value. Wavelength calibration was then applied to the one-dimensional spectra using Xe+Ne+HgAr lamps. This procedure was applied to the spectra of the comet and the stars (both spectrophotometric and solar analogs). We divided the individual spectra of the comet by the spectra of the solar analogs, and the resulting ratio were averaged to get the final reflectance spectra. To create the final, composite reflectance spectrum shown in Fig. 2, we joined the R300B and R300R using the common wavelength interval (6000–7000 Å), and normalized the result to unity at 5500 Å.
A.2 Imaging
Photometric follow-up was carried out on the nights of July 2, 3 and 5, 2025, using one of the two telescopes of the Two-meter Twin Telescope (TTT), located at the Teide Observatory, Canary Islands (Lat. 28∘17′56′′ N, Long. 16∘30′35′′W, Alt. 2361.9 m). This is a 2-m /6 Ritchey-Chrétien telescope that is currently in its commissioning phase. An Andor iKon-L 936 2k2k camera is mounted at the Nasmyth 2 focus, equipped with a back-illuminated 13.5 m pixel-1 BEX2-DD CCD sensor, resulting in a field of view of 7.85′7.85′ and a plate scale of 0.23′′ pixel-1. The object transited the Galactic plane (16, 11°, on July 2), a region with both very high stellar density and significant differential extinction. Given that this extinction strongly affects bluer bands, we decided to observe with a filter to reduce source blending. On the nights of July 3 and 5, a dithering pattern was used to avoid the potential effects of gradients across the sensor caused by the presence of the Moon and high airmass on the photometric quality of the data. A brief description of the image processing methods is included in Appendices B and D.
Appendix B Deep imaging data reduction
Each raw frame was first corrected for a master bias (median‐stack of 21 bias frames), saturated pixels were masked, and the flat-field correction was performed using twilight flats. A comprehensive sky mask using aggresive Gnuastro’s NoiseChisel (Akhlaghi & Ichikawa, 2015) segmentation was applied, and the sky level was measured via a sigma‐clipped median of the unmasked pixels and subtracted, ensuring removal of background without oversubtracting the faint coma.
An initial WCS solution was obtained using Astrometry.net (Lang et al., 2010) against Gaia DR3, refined through three SCAMP (Bertin, 2006) iterations. Each image was then cropped to a 3’3’ region centered on the comet, using coordinates provided by JPL Horizons. To avoid potential astrometric uncertainties, the center was re-calculated from the object’s brightness centroid. Cutout images were reastrometrized and resampled onto a common grid with SWarp (Lanczos3 method, Bertin et al., 2002).
After calibration and sky subtraction, we examined all 326 frames for residual contamination (e.g., faint field stars or tracking artifacts) and excluded 99 images with any potential contamination to the comet signal. The remaining 227 frames, totaling 3.15 hours of integration, were combined using an inverse variance–weighted mean, where each image’s weight was set by the reciprocal of its background variance. In this combination, Noisechisel was used for masking every source but the comet and a 3 clipping was applied to reject outliers.
Appendix C Gas production rate computation
To study the presence of typical cometary emission lines/bands of CN, C2, C3 or NH2 usually detected in the visible range in active comets (see, e.g., Cochran et al. 2012), we first obtained the continuum-subtracted spectrum of 3I/ATLAS (see Fig.5) using a spectrum of the Sun downloaded from the CALSPEC compilation (Bohlin et al., 2014) as we did in de León et al. (2020). We first noticed that there is no evidence of any gas emission in our spectrum. As the CN (0–0) emission band at 3880 Å use to be the most conspicuous one in comets, we computed the upper limit of the detectable CN flux within the SNR of our spectrum in the 3830–3905 Å range (see Fig.5 inset plot) and derived the upper limit of the CN production rate (Q(CN)lim) as follows: i) we first computed the of the flux in the continuum-subtracted spectrum in the 3830–3905 Å spectral range and obtained a value of ; ii) assuming a detection limit of 3- as ; iii) we finally converted into column density using the factor from Schleicher (2010), scaled to both the heliocentric distance and the velocity of the comet. To compute the gas production rate, we assumed the Haser modeling with the outflow velocity scaled with (= km s-1), customary values for the daughter velocity =1 km s-1 , and scale lengths given in A’Hearn et al. (1995). For the corresponding set of parameters in the Haser modeling, we produced theoretical column density profiles for CN varying the production rate until the best match between observations and theoretical predictions was achieved. We obtained mol s-1 .
Appendix D Period-finding method
All raw TTT frames were first corrected for bias and twilight flat‐field variations using standard procedures. Point‐source photometry was then performed with a GPU‐accelerated custom pipeline that employs PSF‐matched convolutions for source detection and adaptive, S/N‐driven aperture photometry with PSF‐based aperture corrections. For each exposure, we matched field stars to the Pan-STARRS PS1-DR1 catalog (Chambers et al., 2016) to derive a photometric zero-point, which was used to calibrate the comet and control-star magnitudes in the resulting time series. This pipeline has been validated in previous works (Popescu et al., 2025) and will be released as open-source software (Alarcon et al., in prep.).
To determine the rotation period of 3I/ATLAS from the TTT time‐series photometry, we began by computing a generalized Lomb–Scargle periodogram (Press & Rybicki, 1989). This method is well suited for unevenly sampled data and quantifies the significance of sinusoidal signals by fitting 2-order sine curves at trial frequencies while properly normalizing for observational errors. Applying it to the combined ‐band light curve revealed a peak at h. To corroborate this result without assuming purely sinusoidal variability, we employed the Phase Dispersion Minimization (PDM) technique, which searches for the period that minimizes the variance of data points within phase bins. Our PDM statistic exhibited a global minimum at a similar period identified by the Lomb–Scargle analysis. We then fitted a Gaussian function to the PDM curve in the vicinity of this minimum (see Fig. 6), yielding a refined rotation period of h. The quoted uncertainty corresponds to the standard deviation of the Gaussian fit and encompasses both measurement noise and sampling effects. We adopt this value as our best estimate in the main text.
To verify that the observed periodic modulation originates from the comet’s nucleus and coma rather than observational artifacts, we investigated potential correlations between measured flux and atmospheric seeing, as well as any drifts in our photometric zero point. In addition, we extracted light curves for several field stars of similar brightness; the bottom panel of Fig. 3 presents a representative non‐variable control star, which shows scatter an order of magnitude smaller than the amplitude of the comet’s modulation. These checks confirm that our period determination is robust against changes in seeing and calibration, and that the periodic signal reflects the intrinsic rotation of 3I/ATLAS.
Appendix E Orbit data
The orbital solution in Table 2 is currently based on 484 observations with an observational time span of 53 days, it is referred to epoch JD 2460860.5 TDB, which is the origin of time in the calculations. It was retrieved from Jet Propulsion Laboratory’s (JPL) Small-Body Database (SBDB)222https://ssd.jpl.nasa.gov/tools/sbdb_lookup.html#/ provided by the Solar System Dynamics Group (SSDG, Giorgini 2011, 2015).333https://ssd.jpl.nasa.gov/ Its current path is hyperbolic with a statistical significance (barycentric) above 645. No known processes, close encounters with known planets included, can explain such a large eccentricity if we were to assume that 3I/ATLAS has an origin in the solar system; an extrasolar origin emerges as the single other alternative.
| Orbital parameter | Heliocentric | Barycentric | |
| Perihelion distance, (au) | = | 1.35800.0011 | 1.3633 |
| Eccentricity, | = | 6.1510.008 | 6.163 |
| Inclination, (°) | = | 175.11370.0006 | 175.1276 |
| Longitude of the ascending node, (°) | = | 322.1700.011 | 322.259 |
| Argument of perihelion, (°) | = | 127.9950.011 | 128.073 |
| Mean anomaly, (°) | = | -855.30.9 | -854.7 |
Appendix F Gaia DR3 sample
An origin in the Galactic thin disk is supported by the fact that several stars in Gaia third data release (DR3) with properties consistent with thin disk membership have kinematic signatures compatible with that of 3I/ATLAS. Gaia DR3 provides estimates of and [Fe/H] for some of the relevant stars and the values are consistent with those of Solar-like stars of Solar-like metallicity as shown in Fig. 7. One star (source 4091535895147524608) with 3I/ATLAS-like kinematics has = 18:34:31.4, =21:39:30, =4580.2 K, [Fe/H]=0.20, a heliocentric distance of 2232.16 pc, and heliocentric Galactic velocity km s-1. Unfortunately, the uncertainties in the values of the radial velocity of the relevant stars in Gaia DR3 are significant and this leads to errors in the components of the velocity which are much larger than those computed for 3I/ATLAS.